# The number of cultural traits, evolving genealogies, and the descendant process

### v1

@@ 3.3.1 Distributions of tree length, $L_{n}$
Figure 10: Histogram of $\log_{10}\ell_{1}$ (green) and $\log_{10}\ell_{N}$ (blue), obtained from $10^{4}$ simulation runs of the untyped ALG; the overlap area of the two histograms appears blue-green. In each run, the ancestors were traced until they disappeared. Top: $s=0.9$. Bottom: $s=1.3$. $N=100,u=1$.
3.3.2 Metastability and coalescence in the ALG
@@
Still, we are at the descriptive level. For a true understanding of the behaviour, we must consider the dynamics of the ALG. The behaviour of the line-counting process $(Y_{n}(\tau))_{\tau\geq 0}$ is very well studied, since, as mentioned above, it is, at the same time, the stochastic logistic process of the SIS model in epidemiology. Recall that $(Y_{n}(\tau))_{\tau\geq 0}$ dies out with probability 1; but it is decisive what happens before extinction. In the subcritical case, $(Y_{n}(\tau))_{\tau\geq 0}$ dies out quickly almost surely. In contrast, in the supercritical regime, it is well known (Andersson and Djehiche, 1998; Foxall, 2021) to have a metastable state around $N\bar{\xi}$, where $\bar{\xi}$ is the stable equilibrium of the differential equation (14). Note that $\lambda_{\lfloor N\bar{\xi}\rfloor}\approx\mu_{\lfloor N\bar{\xi}\rfloor}$; also recall from Section 2.6 that the qualitatively different behaviour in the sub- and supercritical regimes reflects the transcritical bifurcation of the equilibria of the differential equation.
@@
From now on, we mainly focus on the supercritical case in order to understand the sawtooth behaviour for $s>u$. For $(Y_{1}(t))_{\tau\geq 0}$ in the supercritical case, there is the following dichotomy: with probability $u/s$, $(Y_{1}(\tau))_{\tau\geq 0}$ goes extinct quickly; otherwise, it grows to reach the metastable state (say $\lfloor N\bar{\xi}\rfloor$ for definiteness) in a short time. $(Y_{N}(\tau))_{\tau\geq 0}$ always moves to $\lfloor N\bar{\xi}\rfloor$ quickly. By a crude approximation, the expected first-passage times from 1 to $\lfloor N\bar{\xi}\rfloor$ (conditional on non-extinction before reaching $\lfloor N\bar{\xi}\rfloor$), and from $N$ to $\lfloor N\bar{\xi}\rfloor$, are both bounded by $N^{2}/u$ (Andersson and Djehiche, 1998, Proof of Lemma 3) [^4] The factor of $1/u$ comes from the fact that Andersson and Djehiche (1998) work with $u=1$.. (According to Andersson and Djehiche (1998), it can be shown with the help of refined arguments that the bound is actually $\mathcal{O}(\log(N))$.)
@@
But the line-counting process alone does not suffice to understand what is really going on; rather, one has to consider the dynamics of the ALG, more precisely, the processes $(\Lambda_{\alpha}(\tau))_{\tau\geq 0}$, $\alpha\in[N]$, and $(\Lambda_{[N]}(\tau))_{\tau\geq 0}$, simulated according to (6b). In line with the dichotomy described above, we observe in Figure 11 that the ancestors of a single individual either die out quickly or grow to a metastable number and survive for a long time; likewise, the number of ancestors of the entire population reaches the metastable regime quickly. But we also see that, soon after both $(\Lambda_{\alpha}(\tau))_{\tau\geq 0}$ and $(\Lambda_{[N]}(\tau))_{\tau\geq 0}$ have reached a metastable size, they become identical. This is simply due to the fact that $\Lambda_{\alpha}(\tau)\subseteq\Lambda_{[N]}(\tau)$ for all $\tau\geq 0$, and the sizes of the two sets perform fluctuations around $N\bar{\xi}$ of order $\mathcal{O}(1/\sqrt{N})$, in the course of which they finally meet and hence become identical. From this point onwards, we have $\Lambda_{\alpha}(\tau)=\Lambda_{[N]}(\tau)$ due to (7).
The coalescence events [^5] Note that these coalescence events are different from those in the ASG, where coalescence means that two lines merge into one. responsible for this happen whenever, for some $\alpha$, a line in $\Lambda_{\alpha}(\tau)$ learns from a line in $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$: then, a parent from $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$ is added to $\Lambda_{\alpha}(\tau)$ and hence removed from $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$. A transition in the reverse direction (that is, an addition of elements to $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$) is not possible: when a line in $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$ learns from a line in $\Lambda_{\alpha}(\tau)$, neither $\Lambda_{\alpha}(\tau)$ nor $\Lambda_{[N]}(\tau)$ will change.
Figure 11: Simulation of the untyped ALG. $\lvert(\Lambda_{[N]}(\tau))_{\tau\geq 0}\rvert$, the number of ancestors of the total population over time (thick grey line), and the $\lvert(\Lambda_{\alpha}(\tau))_{\tau\geq 0}\rvert$, the sizes of the ancestries of each individual $\alpha$ (100 thin black lines for $\alpha\in[N]$) are shown from backward time $\tau$=0 to 40, all based on the same realisation of $(\Lambda(\tau))_{\tau\geq 0}$. By definition, the thick grey line starts at $N$, and each black line starts at 1. When a black line merges into the grey line (26 occurrences among 100, which is close to the probability of $0.231-u/s=$ 0.23), their ancestor sets become identical and behave identically from then on. $N=100,s=1.3,u=1$.
@@
All this carries over to the traits and explains the stationary behaviour observed in Section 3.2.4. In the subcritical case, the ancestry of all individuals dies out quickly without much of a chance to coalesce, so they only acquire few traits in a more or less independent way; as a consequence, $C_{1}$ and $C_{N}$ both remain small with little dependence between them. This changes in the supercritical case. If the ancestry of a given individual dies out quickly, it will again acquire no or only a few traits; but if it has old ancestors, it will accumulate numerous traits. Moreover, due to (relatively fast) coalescence, the set of old ancestors, and thus the set of traits, is largely shared with the entire population. This explains the bimodal behaviour, that is, $L_{1}$ and $C_{1}$ are either very small or close to $L_{N}$ and $C_{N}$, respectively. Put differently, cultural traits that are old but not too old to have gone extinct are carried by all knowledgeable individuals. So, for old traits that still exist in a population, there are two main types of individuals: those who know none of them and those who know all of them. Altogether, the trait diversity between individuals is low.
@@ 3.3.3 Evolving genealogies
So far, we have understood the stationary behaviour on the basis of $C_{n}\sim\text{Poi}(\mu L_{n})$. It remains to understand the dynamics, that is, the sawtooth behaviour in the supercritical case. To this end, and once more on the basis of the Poissonian relationship between the number of traits and the tree length, we now consider the tree length as a function of forward time $t$ in a stationary version of the forward process, coupled across all times, and define it as
$L_{n}^{t}\mathrel{\mathop{:}}=\int_{0}^{t}\lvert\Lambda_{[n]}^{t}(\tau)\rvert\mathrm{d}\tau=\int_{0}^{T^{t}}\lvert\Lambda_{[n]}^{t}(\tau)\rvert\mathrm{d}\tau,\quad n\in[N],$ (32)
where $\Lambda_{[n]}^{t}(\tau)$ is the set of ancestors at backward time $\tau$ of individuals $1,2,\ldots,n$ at forward time $t$, see Figure 12 for an illustration; and $T^{t}$ is the extinction time of $(\Lambda^{t}_{[N]}(\tau))_{\tau\geq 0}$, which we take to be equal to $t$ if the process has not died out until backward time $t$. We clearly have
@@
As before, the set $[n]$ is representative of any sample of size $n$ due to exchangeability. Let us also note that the $T^{t}$ are all identically distributed in the same way as $T$ of (31), but they are, in general, not independent; and likewise for $L_{n}^{t}$ and $\Lambda_{[n]}^{t}$. All this leads us to the concept of evolving genealogies, as previously studied in the context of the neutral coalescent process (that is, without selection) by Pfaffelhuber et al. (2011). The time course of $(C_{n}(t))_{t\geq 0}$ will follow from there via $C_{n}(t)\sim\textrm{Poi}(\mu L_{n}^{t})$.
Figure 12: The total branch length $L_{N}^{t}$ of the entire population as a function of forward time $t$. The realisation of the forward model, as given by the positions and types of all events, is identical in both panels. The left (right) panel shows a case with $0<t<t_{\rm{max}}$ ($t=t_{\rm{max}}$). The sum of the lengths of the red line segments is $L_{N}^{t}$.
@@
We have extracted the evolving genealogies from a single long simulation run of the forward model in the supercritical case, where we have stored all events (including the information about who died and who learned from whom) and the times at which they happen. This way, we obtained the realisations of $(L^{t}_{N})_{t\geq 0}$ and $(L^{t}_{1})_{t\geq 0}$ coupled across all times, as shown in Figure 13.
@@
Figure 13: The tree length (red) and the number of traits (blue) as functions of forward time $t$ in the same realisation of a forward simulation. Top: $\mu(\ell_{N}^{t})_{t\geq 0}$ and $(c_{N}(t))_{t\geq 0}$. Bottom: $\mu(\ell_{1}^{t})_{t\geq 0}$ and $(c_{1}(t))_{t\geq 0}$. In the bottom panel, they both oscillate quickly between near-0 and close to $\mu\ell_{N}^{t}$, so only points (instead of lines) are shown. $N=100,s=1.3,\mu=0.1,u=1$.
@@
Figure 14: Relationship between (from top to bottom) the total tree length, the set of ancestors, the time to the oldest ancestor, and the descendant process, all extracted from the same forward simulation, and all as a function of forward time $t$. In detail: the top panel shows a realisation $(\ell_{N}^{t})_{t\geq 0}$ of $(L_{N}^{t})_{t\geq 0}$ with a mass extinction at $t\approx 9688$. The second panel displays the number of ancestors at $t\leq t_{i}$ ($i=1,2,3,4)$ of samples taken at forward times $t_{1}=9681,t_{2}=9706,t_{3}=9748,t_{4}=9789$. More precisely, the thick coloured lines are realisations of the $\lvert\Lambda^{t_{i}}_{[N]}(t_{i}-t)\rvert_{t\leq t_{i}}$, whereas the thin grey lines represent realisations of $\lvert\Lambda^{t_{i}}_{\alpha}(t_{i}-t)\rvert_{t\leq t_{i}}$ for all $\alpha\in[N]$; these lines get darker when more lines overlap. The third panel shows the realisation $(s_{N}(t))_{t\geq 0}$ of $(S_{N}^{t})_{t\geq 0}$, the time to the oldest ancestor. The bottom panel displays the number of descendants at time $t\geq t_{i}$ of samples taken at the forward times $t_{i}$. In analogy with the second panel, the thick coloured lines and the thin grey lines are realisations of the $\lvert\Gamma^{t_{i}}_{[N]}(t-t_{i})\rvert_{t\geq t_{i}}$ and the $\lvert\Gamma^{t_{i}}_{\alpha}(t-t_{i})\rvert_{t\geq t_{i}}$, respectively. Looking backward (ancestral process) or forward (descendant process) in time, the thick coloured lines (along with the thin black ones that have coalesced into them) merge and fluctuate around their metastable values for a long time before going extinct. $N=100,u=1,s=1.3.$

### v2

@@ 3.2 Marginal and joint distributions of $L_{1}$ and $L_{N}$, and of $C_{1}$ and $C_{N}$
Figure 6: Histograms of $\log_{10}\ell_{1}$ (green) and $\log_{10}\ell_{N}$ (blue), obtained from $10^{4}$ simulation runs of the untyped ALG; the overlap area of the two histograms appears blue-green. In each run, the ancestors were traced until they disappeared. Top: $s=0.9$; bottom: $s=1.3$. $N=100,u=1$.
In the subcritical case, the histograms simply tell us that single individuals tend to have short ancestries, whereas the ancestries of the entire population are much longer (note the log scale). In the supercritical case, the values of $\ell_{N}$ become much larger, and the distribution of $\ell_{1}$ is bimodal: it has a large peak for small values, where most of the mass is located, and a second smaller peak where the distribution of $\ell_{N}$ is located. So most individuals have short genealogies, but occasionally they become as long as the genealogy of the entire population.
Recalling now that $C_{n}\sim\text{Poi}(\mu L_{n})$, we turn the histograms of $\ell_{1}$ and $\ell_{N}$ into those of $c_{1}$ and $c_{N}$, where $c_{n}$ is a realisation of $C_{n}$, see Figure 7. A qualitative difference appears only for the left peaks: the left peaks of the histograms of $\ell_{1}$ translate into the left peaks of the histograms of $c_{1}$ in that short genealogies mean high chances for no or only a very small number of innovation events, resulting in a sharp peak of (nearly) ignorant individuals. So the histograms show that, in the supercritical case, most individuals are (almost) ignorant, but some of them carry nearly all the knowledge of the entire population.
This becomes even clearer when we consider the joint distribution of $\ell_{1}$ and $\ell_{N}$ in heat maps, and likewise for $c_{1}$ and $c_{N}$ (Fig. 8). In the subcritical case, the correlation between $\ell_{1}$ and $\ell_{N}$ is altogether weak; $\ell_{1}$ is bimodal in the rare case that $\ell_{N}$ is relatively large. A similar tendency exists between $\ell_{2}$ and $\ell_{N}$ (not shown). In contrast, the correlation is very strong in the supercritical case. When $\ell_{N}$ is not too small, the distribution of $\ell_{1}$ is bimodal, and there are two extreme types of individuals: those with very short genealogies and those whose genealogies are nearly as long as those of the entire population; there are only very few intermediate cases. An analogous observation applies to $\ell_{2}$ and $\ell_{N}$ (not shown).
Unsurprisingly, a very similar picture emerges in the joint distribution of $c_{1}$ and $c_{N}$. The only difference is some additional noise caused by the Poisson random variable translating $\ell_{n}$ into $c_{n}$.
[…]
Figure 8: Heat maps showing the joint distributions of $\ell_{1}$ and $\ell_{N}$ (blue) and of $c_{1}$ and $c_{N}$ (orange), where the values of $\ell_{1},\ell_{N},c_{1}$, and $c_{N}$ are those from Figures 6 and 7. Top: $s=0.9$, bottom: $s=1.3$. $N=100,u=1$.
3.3 Metastability and merging in the ALG
@@
Still, we are at the descriptive level. For an understanding of the behaviour, we must consider the dynamics of the ALG. As mentioned above, the line-counting process $(Y_{n}(\tau))_{\tau\geq 0}$ dies out with probability 1; but it is decisive what happens before extinction. Let us briefly recapitulate the essentials following Andersson and Djehiche (1998) and Foxall (2021). In the subcritical case, $(Y_{n}(\tau))_{\tau\geq 0}$ dies out quickly almost surely. In contrast, in the supercritical regime, it has a metastable state around $N\bar{\xi}$, where $\bar{\xi}$ is the stable equilibrium of the differential equation (14). Note that $\lambda_{\lfloor N\bar{\xi}\rfloor}\approx\mu_{\lfloor N\bar{\xi}\rfloor}$; also recall from Section 2.6 that the qualitatively different behaviour in the sub- and supercritical regimes reflects the transcritical bifurcation of the equilibria of the differential equation.
@@
For $(Y_{1}(t))_{\tau\geq 0}$ in the supercritical case, there is the following dichotomy: with probability $u/s$, $(Y_{1}(\tau))_{\tau\geq 0}$ goes extinct quickly; otherwise, it grows to reach the metastable state (say $\lfloor N\bar{\xi}\rfloor$ for definiteness) in a short time. $(Y_{N}(\tau))_{\tau\geq 0}$ always moves to $\lfloor N\bar{\xi}\rfloor$ quickly. By a crude approximation, the expected first-passage times from 1 to $\lfloor N\bar{\xi}\rfloor$ (conditional on non-extinction before reaching $\lfloor N\bar{\xi}\rfloor$), and from $N$ to $\lfloor N\bar{\xi}\rfloor$, are both bounded by $N^{2}/u$ (Andersson and Djehiche, 1998, Proof of Lemma 3) [^4] The factor of $1/u$ comes from the fact that Andersson and Djehiche (1998) work with $u=1$.. (According to Andersson and Djehiche (1998), it can be shown with the help of refined arguments that the bound is actually $\mathcal{O}(\log(N))$.)
@@
But the line-counting process alone does not suffice to understand what is really going on; rather, one has to consider the dynamics of the ALG, more precisely, the processes $(\Lambda_{\alpha}(\tau))_{\tau\geq 0}$ for $\alpha\in[N]$ and $(\Lambda_{[N]}(\tau))_{\tau\geq 0}$ of (6b). In line with the dichotomy described above, the simulations in Figure 9 show that the ancestors of a single individual either die out quickly or grow to a metastable number and survive for a long time; likewise, the number of ancestors of the entire population reaches the metastable regime quickly. But we also see that, soon after both $(\Lambda_{\alpha}(\tau))_{\tau\geq 0}$ and $(\Lambda_{[N]}(\tau))_{\tau\geq 0}$ have reached a metastable size, they become identical. This is due to the fact that $\Lambda_{\alpha}(\tau)\subseteq\Lambda_{[N]}(\tau)$ for all $\tau\geq 0$, and the sizes of the two sets perform fluctuations around $N\bar{\xi}$ of order $\mathcal{O}(1/\sqrt{N})$, in the course of which they finally meet and hence become identical. From this point onwards, we have $\Lambda_{\alpha}(\tau)=\Lambda_{[N]}(\tau)$ due to (7).
The merging events [^5] Note that these merging events differ from the coalescence events in the ASG, where two lines unite into one. responsible for this happen whenever, for some $\alpha$, a line in $\Lambda_{\alpha}(\tau)$ learns from a line in $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$: then, a parent from $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$ is added to $\Lambda_{\alpha}(\tau)$ and hence removed from $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$. A transition in the reverse direction (that is, an addition of elements to $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$) is not possible: when a line in $\Lambda_{[N]}(\tau)\setminus\Lambda_{\alpha}(\tau)$ learns from a line in $\Lambda_{\alpha}(\tau)$, neither $\Lambda_{\alpha}(\tau)$ nor $\Lambda_{[N]}(\tau)$ will change.
Figure 9: Simulation of the untyped ALG. $\lvert(\Lambda_{[N]}(\tau))_{\tau\geq 0}\rvert$, the number of ancestors of the total population over time (thick grey line), and the $\lvert(\Lambda_{\alpha}(\tau))_{\tau\geq 0}\rvert$, the sizes of the ancestries of each individual $\alpha$ (100 thin black lines for $\alpha\in[N]$) are shown from backward time $\tau$=0 to $\tau=40$, all based on the same realisation of $(\Lambda(\tau))_{\tau\geq 0}$. By definition, the thick grey line starts at $N$, and each thin black line starts at 1. When a black line merges into the grey line (26 occurrences among 100, which is close to the probability of 1u/s=0.231-u/s=0.23), the corresponding sets become identical and behave identically from then on. $N=100,s=1.3,u=1$.
@@
All this explains the stationary behaviour illustrated in Figure 8. In the subcritical case, the ancestries of all individuals die out quickly without much of a chance to merge, so most of them remain short and more or less independent. This changes in the supercritical case. If the ancestry of a given individual dies out quickly, it will again remain short and independent of $L_{N}$; this corresponds to the concentration of points close to the horizontal axis in the corresponding heat map. But if the number of ancestors moves to the metastable state, it will stay there for a long time, so the individual has old ancestors. Moreover, due to (relatively fast) merging, the set of old ancestors, and thus the set of traits, is largely shared with the entire population. This corresponds to the cloud of points close to the diagonal. As a consequence, cultural traits that are old but not too old to have gone extinct are carried by all knowledgeable individuals. So, for old traits that still exist in a population, there are two main types of individuals: those who know none of them and those who know all of them. Altogether, the trait diversity between individuals is low.
Let us mention that a fast merging of ancestral sets similar to that observed above was described, and actually proved, for the minimal-load ASGs in a model of Muller’s ratchet with tournament selection (González Casanova et al., 2023, Sec. 7). Like the ALG, the minimal-load ASG experiences branching and pruning events; but it additionally has coalescence events, which are crucial for the proof, so the latter does not carry over to our situation.
3.4 Forward dynamics and evolving genealogies
So far, we have considered the stationary distributions of $L_{n}$ and $C_{n}$. Let us now return to the dynamics $(L_{n}^{t})_{t\geq 0}$ and $(C_{n}(t))_{t\geq 0}$; recall that $(L_{n}^{t})_{t\geq 0}$ is the length of the evolving genealogy, and the upper index refers to the time where we start looking back.
3.4.1 Forward simulations
We carried out individual-based simulations of the forward process $(\Phi(t))_{t\in[0,t_{\rm{max}}]}$ described in Section 2.2. We start with $k_{i}(0)=\varnothing$ for every $i\in[N]$ and compute realisations $c_{n}(t)$ of $C_{n}(t)$ at $t=0,1,...,t_{\rm{max}}$. We take $t_{\rm{max}}$ large enough so that the effect of the initial state can be neglected and denote by $t_{\rm{max}}]_{0}$ the set $\{0,1,\ldots,t_{\rm{max}}\}$.
@@ 3.4.2 Evolving genealogies
We are now ready to consider $L_{n}^{t}$, the length of the genealogy as a function of the forward time $t$ where we start looking back, as defined in (10); and we understand it as the process $(L_{n}^{t})_{t\geq 0}$ coupled across all times via the graphical representation, as illustrated in Figure 10. The same applies, of course, to the $(\Lambda^{t}_{[n]}(\tau))_{\tau\geq 0}$. We clearly have
@@
As before, the set $[n]$ is representative of any sample of size $n$ due to exchangeability. The coupling across times extends to the extinction times $T^{t}$ of $(\Lambda^{t}_{[N]}(\tau))_{\tau\geq 0}$; the $T^{t}$ are all identically distributed in the same way as $T$ of (29), but they are, in general, not independent.
@@
All this leads us to the concept of evolving genealogies, as previously studied in the context of Kingman’s coalescent process (that is, with coalescence, but without branching events) by Pfaffelhuber and Wakolbinger (2006) and Pfaffelhuber et al. (2011). In practice, we have extracted the evolving genealogies from a single long simulation run of the forward model, where we have stored all events (including the information about who died and who learned from whom) and the times at which they happen. This way, we obtained realisations of $(\Lambda^{t}_{n})_{t\geq 0}$ and $(L^{t}_{n})_{t\geq 0}$ coupled across all times. We will now discuss them together with the realisations of $(C_{n}(t))_{t\geq 0}$ extracted directly from the forward simulation.
Figure 10: The total length $L_{N}^{t}$ of the genealogy of the entire population as a function of forward time $t$. The realisation of the graphical representation, as given by the positions and kinds of all events, is identical in both panels. The top (bottom) panel shows a case with $0<t<t_{\rm{max}}$ (with $t=t_{\rm{max}}$). The sum of the lengths of the red line segments is $L_{N}^{t}$.
3.4.3 Forward dynamics: length of genealogy and number of traits
Long time series $(\mu\ell_{N}(t))_{t\geq 0}$, $(c_{N}(t))_{t\geq 0}$, $(\mu\ell_{1}(t))_{t\geq 0}$, and $(c_{1}(t))_{t\geq 0}$ are shown in Figures 11 and 12 for the subcritical and the supercritical case, respectively. More detail is to be seen in Figure 13, which shows short cutouts of $(\ell_{N}(t))_{t\geq 0}$ and $(\ell_{1}(t))_{t\geq 0}$ in both cases. In the subcritical case, both $(\ell_{1}(t))_{t\geq 0}$ and $(\ell_{N}(t))_{t\geq 0}$ fluctuate moderately around some constant value in a more or less symmetric manner, and likewise for the traits. The number of traits seems to be largely independent of the length of the genealogies.
In contrast, in the supercritical case, the time series $(\mu\ell_{N}^{t})_{t\geq 0}$ and $(c_{N}(t))_{t\geq 0}$ are very close in our simulations, and likewise for $(\mu\ell_{1}^{t})_{t\geq 0}$ and $(c_{1}(t))_{t\geq 0}$. Before we turn to the conspicuous shape of the supercritical dynamics, let us comment on the different degrees of coupling of $(\ell_{n}^{t})_{t\geq 0}$ and $(c_{n}(t))_{t\geq 0}$ (for $n\in\{1,N\}$) in the sub- and supercritical cases. It reflects the variance components of $C_{n}$ in Section 3.1.3 and Figure 4: recall that the variance due to the innovation process is an appreciable proportion of $\mathbb{V}[C_{n}]$ in the subcritical case, but is negligible for the supercritical case, so $\mathbb{V}[C_{n}]$ is then largely governed by the variability of the length of the genealogy. This explains that $(\ell_{n}^{t})_{t\geq 0}$ and $(c_{n}(t))_{t\geq 0}$ fluctuate more or less independently for $s<1$, but are closely coupled for $s>1$.
As to the dynamics in the supercritical case, both $(\ell_{N}(t))_{t\in[t\geq 0}$ and $(c_{N}(t))_{t\geq 0}$ display a characteristic sawtooth behaviour: they increase more or less linearly (at roughly constant slope in every sawtooth) for some period, followed by a rapid collapse to a small value at a random time. We will find the reason for this in the next section.
The dynamics of $(\ell_{1}(t))_{t\geq 0}$ and $(c_{1}(t))_{t\geq 0}$ in the supercritical case confirm the dichotomy observed in Section 3.3: individual ancestries quickly either die out or move to the metastable state, so they are either very short or nearly as long as the genealogy of the entire population; as a consequence, individuals know ‘nearly nothing’ or ‘nearly everything’. The dynamics yields the additional insight that the role of any given individual (short or long genealogy) changes quickly, typically many times within a given sawtooth, see the lower panel of Figure 12; this happens when an individual with a short history is attached to an individual with a long history via a learning arrow; or, the other way around, if an individual with a long history dies.
As a consistency check, we have also produced the histograms of $(\ell_{1}(t))_{t\in[t_{\text{max}}]_{0}}$, $(c_{1}(t))_{t\in[t_{\text{max}}]_{0}}$, $(\ell_{N}(t))_{t\in[t_{\text{max}}]_{0}}$, and $(c_{N}(t))_{t\in[t_{\text{max}}]_{0}}$ resulting from our time series, and they are practically indistinguishable from the histograms in Figures 6 and 7, as it should be by stationarity and ergodicity.
@@ 3.4.3 Forward dynamics: length of genealogy and number of traits
Figure 11: The innovation rate times the length of the genealogy (red) and the number of traits (blue) as functions of forward time $t$ in the same realisation of a forward simulation. Top: $\mu(\ell_{N}^{t})_{t\geq 0}$ and $(c_{N}(t))_{t\geq 0}$. Bottom: $\mu(\ell_{1}^{t})_{t\geq 0}$ and $(c_{1}(t))_{t\geq 0}$. In the bottom panel, both time series oscillate quickly between near-0 and close to $\mu\ell_{N}^{t}$, so we did not interpolate between the points. $N=100,s=0.9,\mu=0.1,u=1$.
Figure 12: The innovation rate times the total length of the genealogy (red) and the number of traits (blue) as functions of forward time $t$ in the same realisation of a forward simulation. Top: $\mu(\ell_{N}^{t})_{t\geq 0}$ and $(c_{N}(t))_{t\geq 0}$. To make the blue curve visible ‘below’ the red, we did not interpolate between the red points. Bottom: $\mu(\ell_{1}^{t})_{t\geq 0}$ and $(c_{1}(t))_{t\geq 0}$. Here, both time series oscillate quickly between near-0 and close to $\mu\ell_{N}^{t}$, so we did not interpolate between any points. $N=100,s=1.3,\mu=0.1,u=1$.
Figure 13: Time series $\big(\ell_{1}(t)\big)_{t\geq 0}$ (green) and $\big(\ell_{N}(t)\big)_{t\geq 0}$ (black) for the last 400 time units in a run with $t_{\rm{max}}=10^{4}$ in the subcritical case (upper panel, $s=0.9$) and the supercritical case (lower panel, $s=1.3$). $N=100,\mu=0.1,u=1$.
@@
Figure 14: Relationship between (from top to bottom) the total length of the genealogy, the set of ancestors, the time to the oldest ancestor, and the descendant process, all extracted from the same forward simulation, and all as a function of forward time $t$. In detail: the top panel shows a realisation $(\ell_{N}^{t})_{t\geq 0}$ of $(L_{N}^{t})_{t\geq 0}$ with a mass extinction (that is, an abrupt collapse to near zero) at $t\approx 9688$. The second panel displays the number of ancestors at $t\leq t_{i}$ ($i=1,2,3,4)$ of samples taken at forward times $t_{1}=9681,t_{2}=9706,t_{3}=9748$, and $t_{4}=9789$. More precisely, the thick coloured lines are realisations of the $\lvert\Lambda^{t_{i}}_{[N]}(t_{i}-t)\rvert_{t\in[t_{i}]_{0}}$, whereas the thin grey lines represent realisations of $\lvert\Lambda^{t_{i}}_{\alpha}(t_{i}-t)\rvert_{t\in[t_{i}]_{0}}$ for all $\alpha\in[N]$; these lines get darker when more lines overlap. The third panel shows the realisation $(s_{N}(t))_{t\geq 0}$ of $(S_{N}^{t})_{t\geq 0}$, the time to the oldest ancestor. The bottom panel displays the number of descendants at time $t\geq t_{i}$ of samples taken at the forward times $t_{i}$. In analogy with the second panel, the thick coloured lines and the thin grey lines are realisations of the $\lvert\Gamma^{t_{i}}_{[N]}(t-t_{i})\rvert_{t\geq t_{i}}$ and the $\lvert\Gamma^{t_{i}}_{\alpha}(t-t_{i})\rvert_{t\geq t_{i}}$, respectively. Looking backward (ancestral process) or forward (descendant process) in time, the thick coloured lines (along with the thin black ones that have merged into them) merge and fluctuate around their metastable value for a long time before going extinct. $N=100,u=1,s=1.3.$
3.5 Understanding the sawtooth: ancestral sets and descendant process